Determine the horizontal force P to be applied to a block of weight of 1800N to hold it in position on a smooth inclined plane, which makes an angle 300 with horizontal reference line.
 
 

Grade:12th pass

2 Answers

Vikas TU
4 years ago
Dear student 
Weight downward = 1800N 
Horizontal force required to keep Block in inclined plane = mgSin30 
mg = 1800N 
Horizontal force = 1800*1/2 
=> 900 N 
Hope this helps 
Joel Sudhakar
4 years ago
dear brother 
here weight acting downward =1800N
the force acting  along the plane to hold the block is =mgSin30=1800*1/2=900
but here we need to calculate the horizontal force ‘P’which is acting on the body 
so, we resolve the horizontal force ‘P’ into x-component and y-component as PCos30 and Psin30 respectively
then Pcos30=mgSin30
P=(1800*1/2)/Cos30=1039.23N